Nitrogen gas is filled in an insulated container. If alpha fraction of moles dissociates without exchange of any energy then the fractional change in its temperature is:
1
Expert's answer
2019-12-09T08:26:09-0500

1)


"U=2.5nRT_1\to T_1=\frac{2E}{5nR}"

2)


"U=(1-\alpha)2.5nRT_2+(\alpha)1.5(2n)RT_2\to T_1=\frac{2E}{(5+\alpha)nR}"

3) The fractional change in its temperature is:


"\frac{\frac{2E}{(5+\alpha)nR}-\frac{2E}{5nR}}{\frac{2E}{5nR}}=\frac{-\alpha}{5+\alpha}"